OpenStudy (anonymous):

n^2+4n-24 factored is

OpenStudy (anonymous):

Simplifying n2 + 4n + -24 Reorder the terms: -24 + 4n + n2 Final result: -24 + 4n + n2

OpenStudy (anonymous):

Correct?

OpenStudy (anonymous):

i guess?

hero (hero):

I don't mean to be condescending, but that's not how you factor

OpenStudy (anonymous):

its mkk

hero (hero):

Proper Steps to factor: 1. Begin with a trinomial of the form \(a^2 + bx + c \) 2. Split the middle term in the following manner: \(ax^2 + (\frac{b}{2} + d)x + (\frac{b}{d} - d)x + c\) d = the root of the polynomial you need to find 3. Factor out the common factor of the first two terms. In this case, ax is common to the first to terms: \(ax(x + (\frac{b}{2a} + d))\) 4. Factor out the common factor of the last two terms. In this case, \(\frac{b}{2} - d\) is common to both: \((\frac{b}{2} - d)(x + \frac{b}{2a} + d)\) 5. Observe after factoring the first two terms and last two terms an expression common to both: \(ax(x + (\frac{b}{2a} + d)) + (\frac{b}{2} - d)(x + \frac{b}{2a} + d)\) 6. Factor out \((x + \frac{b}{2a} + d)\) from both terms: (\(x + \frac{b}{2a} + d)(ax + (\frac{b}{2} - d)\)

hero (hero):

That's of course the general method of factoring